If Φ (r)= 2^r, show that Φ (r+1) =2 Φ(r).

1
Expert's answer
2022-06-06T13:28:56-0400

"\u03a6 (r)= 2^r,"

then "\Phi(r+1)=2^{r+1}=2^r\cdot2^1=2\cdot2^r=2\Phi(r),"

so the statement "\u03a6 (r+1) =2 \u03a6(r)" holds.